# The Cost of Dynamic Reasoning: Demystifying AI Agents and Test-Time Scaling from an AI Infrastructure Perspective

### v1

@@ V Demystifying Test-Time Scaling in AI Agents
In the previous section, we characterized the system-level behavior of AI agents by analyzing their workflow patterns, GPU resource utilization, and serving characteristics. We now shift our focus to the diverse design space of AI agents and examine their test-time scaling behavior to understand the trade-offs between model accuracy and deployment cost. Section V-A first explores the cost-efficiency of various AI agent designs by quantifying how differences in design parameters affect both accuracy and cost-effectiveness. Section V-B then analyzes the test-time scaling behavior of AI agents by comparing multiple scaling strategies and evaluating their impact on accuracy and compute cost. To assess each configuration, we used a benchmark of 50 sample questions and measured the average accuracy and compute cost for each.
@@ V-A Analyzing Cost-Efficiency Across AI Agent Design Spaces
(a) Accuracy
(b) Accuracy per latency
(c) Accuracy per FLOPs
Figure 18: Accuracy and cost-efficiency of AI agent design points. (a) Accuracy vs. end-to-end latency. (b) Accuracy per latency, and (c) Accuracy per FLOPs. (b) and (c) illustrates how efficiently each configuration translates compute cost into task performance.
@@
Figure 18(b) and (c) illustrate the cost-efficiency of various agent configurations. We define cost-efficiency as the ratio of accuracy to cost, where cost is measured either as end-to-end latency (Figure 18(b)) or estimated FLOPs (Figure 18(c)). This metric reflects how effectively each configuration translates compute resources into task accuracy. Across all agents and workloads, we observe a consistent pattern: as computation cost increases, accuracy improves, but with diminishing returns. This highlights the importance of careful system design. Instead of optimizing purely for accuracy, it is critical to identify configurations near the Pareto frontier, where the balance between accuracy and cost is most optimal.
Figure 19: Latency and accuracy trends under iteration budget constraints in ReAct. As the maximum iteration budget increases, both accuracy and average latency initially improve but eventually saturate, while 95th percentile latency continues to grow. Markers indicate the points of maximum accuracy (red diamond) and peak cost-efficiency (blue diamond), as measured by accuracy-to-latency ratio.
Figure 20: End-to-end latency and accuracy trends with varying numbers of few-shot examples in ReAct. With more examples, accuracy initially improves while latency decreases, but both metrics eventually saturate. Markers indicate the configuration with the highest accuracy (red diamond) and the point of peak cost-efficiency (blue diamond), based on normalized accuracy-to-latency ratio.
@@
Key takeaway #10: Design choices in AI agents, such as increasing iteration depth or adding few-shot examples, can improve accuracy, but often show diminishing returns or even regressions beyond a threshold. These improvements also come with significant latency costs. To ensure cost-effective operation, AI agent serving systems should aim to maximize accuracy per unit of compute by selecting configurations that balance quality and responsiveness under real-world constraints.
(a) Reflexion (sequential)
(b) LATS (sequential)
(c) LATS (parallel)
Figure 21: Accuracy–latency trade-offs under sequential and parallel test-time scaling on HotpotQA. Legends denote the scaling level: maximum reflection steps in (a, b), and number of child nodes per expansion in (c).
@@ V-B Test-Time Scaling of AI Agents
Figure 21(a) and (b) show the accuracy–latency trade-offs for Reflexion and LATS under sequential scaling. Both methods improve in accuracy with more reflection steps, but with diminishing returns. For example, in Reflexion, increasing latency from 16.9s to 25.6s yields a 4% accuracy gain. However, achieving the same model accuracy improvement from a later point (56.0s) requires a much larger increase in latency (269.5s), a 31$\times$ higher cost for the same marginal gain.
@@
Figure 22: Accuracy–cost trade-offs under test-time scaling across two model sizes (Llama-3.1-Instruct 8B and 70B) on HotpotQA. (a)–(c) compare Reflexion (top row) and LATS (bottom row) across latency, token usage, and energy consumption. While 70B achieves higher accuracy with fewer steps, the 8B model, especially when paired with parallel scaling, can approach 70B performance with lower energy cost. Each point corresponds to a different level of test-time reasoning.
@@
Figure 22(a) shows that both the 8B and 70B Llama-3.1-Instruct [25, 24] models eventually reach saturation in accuracy, but they differ in how quickly they reach this point. The 70B model achieves high accuracy with relatively low latency, whereas the 8B model requires much longer inference times to reach similar performance. This trend is echoed in Figure 22(b), which plots total token usage. The 8B model consumes significantly more tokens at high-accuracy settings, indicating it needs more reasoning iterations to match the 70B model’s performance. However, as shown in Figure 22(c), the 8B model is substantially more energy-efficient. While the 70B model relies on 8 A100 GPUs, the 8B model runs on just one, resulting in lower total energy consumption per request, even when requiring more reasoning steps to be involved.
Interestingly, the performance gap between models can be partially closed with effective scaling strategies. Reflexion (which uses sequential scaling) shows limited accuracy on the 8B model. But with LATS and parallel scaling, the 8B model achieves near-70B performance by exploring multiple paths and selecting the best one. This shows that test-time strategy can play a compensatory role in low-resource settings.
Key takeaway #11: AI agents scale their reasoning at test time either by deeper sequential thinking or by exploring multiple reasoning paths in parallel, and each strategy imposes distinct trade-offs in latency and resource usage. Sequential scaling favors lower peak resource usage but incurs longer runtime, while parallel scaling accelerates inference at the cost of increased compute and memory demand. System designers should match the appropriate scaling strategy to deployment constraints, potentially combining both for optimal cost-effectiveness.
Accuracy (%) Latency (seconds) Energy (Wh/query) Power @ 71.4 Million Queries/day (Watt) Power @ 13.7 Billion Queries/day (Watt)
8B ShareGPT 4.23 (1×) 0.32 (1×) 1.0 M 182.7 M
Reflexion 38 649.34 (153.7×) 41.53 (130.9×) 123.6 M 23.7 G
LATS 80 380.90 (90.1×) 22.76 (71.7×) 67.7 M 13.0 G
70B ShareGPT 6.40 (1×) 2.55 (1×) 7.6 M 1.5 G
Reflexion 67 720.00 (112.6×) 348.41 (136.5×) 1.0 G 198.9 G
LATS 82 305.67 (47.8×) 158.48 (62.1×) 471.5 M 90.5 G
@@
Figure 23: Growth trend in weekly active users of OpenAI’s ChatGPT. The data shows a marked acceleration in adoption, with the user base surpassing 500 million by April 2025 [39, 40, 31, 36, 41, 35].
VI AI Infrastructure Implications of
Agentic Test-Time Scaling
To assess the system-level impact of agentic test-time scaling, we quantify GPU energy consumption and datacenter-wide power demands of AI agents. Following the methodology in Section V-B, we utilize Reflexion and LATS as representative AI agents employing sequential and parallel scaling, respectively (Table III).
Energy consumption (per query). Using Llama-3.1-Instruct 8B and 70B as backend LLMs, Reflexion consumes 41.53 Wh and 348.41 Wh per request, whereas LATS consumes 22.76 Wh and 158.48 Wh, respectively. By contrast, a conventional single-turn LLM inference based on ShareGPT requires only 0.32 Wh (8B) and 2.55 Wh (70B) per request. These figures correspond to a 62.1$\times$–136.5$\times$ increase in GPU energy per query under agent-based test-time scaling.
Based on recent estimates, ChatGPT serves roughly 500 million weekly active users (WAU) [35] (Figure 23), which translates to about 71.4 million daily active users (DAU). Assuming, conservatively, that each user submits only one agentic query per day, Reflexion’s daily GPU energy footprint would be roughly 2.97 GWh for the 8B model and 24.89 GWh for the 70B model. While our analysis does not consider LLM request batching [66, 19], our estimate is still conservative because (1) it reflects current DAU with only one query per user, even though AI adoption continues to accelerate often with much heavier per-user usage, and (2) it counts only GPU energy, omitting CPU, memory, networking, storage, and cooling overheads.
Even under these modest assumptions, the projected demand rivals the daily electricity consumption of Seattle and its surrounding area (24.8 GWh) [43]. As AI agents become increasingly embedded in everyday applications, their query volume could approach, or exceed, that of traditional search engines. For instance, Google Search processes over 13.7 billion queries per day [12], roughly 192$\times$ the 71.4 million agentic queries assumed above. If this growth in user base and usage persists, AI-infrastructure demand could rise dramatically, potentially exceeding sustainable limits and underscoring the significant challenges posed by test-time scaling.
Datacenter-wide power demands. We now move on to estimating the datacenter-wide power requirements to sustain the aforementioned AI service demands, assuming today’s (ChatGPT’s current 71.4 million queries per day, assuming one query per user) and tomorrow’s (Google search’s 13.7 billion queries per day) AI traffic. The last two columns of Table III translate the per-query energy consumption numbers into datacenter-level power requirements. Under today’s 71.4 million DAU load, single-turn ShareGPT (70B) requires roughly 7.6 MW, well within the tens-of-megawatts envelope typical of modern datacenters [15, 59]. However, assuming similar traffic levels for AI agents, even the lighter 8B-based agents demand 67.7–123.6 MW, comparable to the power draw of a mid-sized U.S. city, while 70B-based agents approach 1 GW, nearly three orders of magnitude higher than the single-turn LLM baseline. Strikingly, this gigawatt-scale power requirement aligns with the announced budget for OpenAI’s multi-gigawatt Stargate facility [32], which is intended to support future AI model deployments. Yet, our analysis suggests that such infrastructure may already be necessary to support agentic systems under today’s traffic levels. Overall, our estimates indicate that even modest user traffic (on the order of tens of millions of queries per day) becomes gigawatt-scale once per-query energy exceeds $\sim$100 Wh, a threshold representative of current agentic workloads.
@@ VI AI Infrastructure Implications of
Sustainability challenges of agentic test-time scaling. Collectively, our findings show that AI agent performance does not scale proportionally with the associated compute, energy, and power costs. Once accuracy saturates, additional test-time scaling yields diminishing returns while imposing substantial system-level burdens. This cost inefficiency is not merely theoretical; it poses concrete constraints on real-world deployments. For instance, OpenAI’s recently introduced Deep Research system [34], designed for complex multi-step reasoning, can take up to 30 minutes per request [33]. To keep infrastructure costs manageable, OpenAI limits usage to 25 runs every 30 days for ChatGPT Plus users and 250 runs for Pro users [33]. These limits highlight the financial and computational challenges of sustaining AI systems that rely heavily on intensive test-time computation.
@@
VII Related Work
AI agent workflows. Recent advances in LLM-based AI agents have introduced diverse workflows that combine language-based reasoning with external tool use. ReAct [64] interleaves reasoning and tool invocation through step-by-step decision making, while Reflexion [47] enhances agent behavior through self-evaluation and feedback. LATS [68] employs a tree-based search strategy to explore multiple reasoning paths in parallel, and LLMCompiler [18] plans tool usage in advance by generating action sequences. Building on these single-agent frameworks, multi-agent systems extend agent capabilities through collaborative task execution. CAMEL [22] implements role-based dialogue among agents, and AutoGen [58] formalizes multi-agent orchestration via structured roles, message passing, and execution control.
While these workflows substantially improve reasoning capabilities and behavioral flexibility, their system-level implications, such as resource usage, latency dynamics, and energy consumption, remain underexplored. To the best of our knowledge, this work is the first to address this gap through a comprehensive system-level analysis of representative AI agent designs, offering new insights into the efficiency and scalability of agentic systems.
@@ VII Related Work
Structured agent interfaces for tool-augmented reasoning. In parallel with behavioral advancements in AI agents, recent efforts have focused on standardizing AI agent APIs and protocols to facilitate broader integration and deployment. OpenAI’s function-calling interface [30] defines a structured mechanism for API invocation, enabling agents to interact with tools in a verifiable and consistent manner. Anthropic’s Model-Context-Protocol (MCP) [3] further formalizes how agents manage context and interact with tools, enabling decoupled reasoning logic and modular deployment across platforms. Google’s Agent-to-Agent (A2A) protocol [13] complements these efforts by specifying a standard for multi-agent communication. A2A enables heterogeneous agents to exchange structured messages, delegate tasks, and coordinate behavior, supporting modular and scalable agentic systems.
Although these efforts focus on unifying the agent programming model, our work takes an orthogonal system-level perspective, uncovering the AI infrastructural challenges posed by agentic workloads under test-time scaling.
System-level optimization of AI agents. Several recent studies have proposed system-level techniques to improve the runtime performance of AI agents by restructuring their execution workflows and addressing latency bottlenecks. LLMCompiler [18], Alto [42], and Teola [49] reduce inference latency by enabling pipelined and parallel execution across reasoning steps. Autellix [23] introduces a queue-aware scheduling policy that minimizes wait time based on latency statistics. AI Metropolis [60] reduces coordination overhead in multi-agent systems via out-of-order execution, while Murakkab [7] improves resource isolation by managing agent workloads at the cloud scheduling layer.
@@
While these works focus on optimizing specific components such as scheduling or execution flow, our study provides a broader characterization of system-level behaviors across diverse agent architectures, helping to understand the resource-efficiency trade-offs of AI agents at scale.
VIII Conclusion

### v2

@@ V Demystifying Test-Time Scaling in AI Agents
We now explore the diverse design space of AI agents and examine their test-time scaling behavior to understand the trade-offs between model accuracy and cost. We evaluated accuracy following the official evaluation protocol of each benchmark. For HotpotQA and MATH, we report exact match accuracy, allowing minor formatting variations (e.g., equivalent mathematical expressions) in MATH. For WebShop, we use the task-specific score defined in the benchmark. For HumanEval, accuracy denotes the proportion of tasks that successfully pass all unit tests. To assess each design point, we used a benchmark of 50 sample questions and measured the average accuracy and the computation cost for each.
@@ V-A Analyzing Cost-Efficiency Across AI Agent Design Spaces
Figure 13: Accuracy and cost-efficiency of AI agent design points. (a) Accuracy vs. latency and (b) Accuracy per latency, illustrating how efficiently each configuration translates cost into task performance.
@@
Figure 13(b) illustrates the cost-efficiency of various agent configurations. We define cost-efficiency as the ratio of accuracy to cost, where cost is measured as end-to-end latency. This metric reflects how effectively each configuration translates compute resources into task accuracy [^3] Using FLOPs as a proxy for cost (“accuracy per FLOP”) yielded similar qualitative conclusions, so we omit those results for brevity.. Across all agents and workloads, we observe a consistent pattern: as computation cost increases, accuracy improves, but with diminishing returns. This underscores the importance of designing AI agent serving systems that find configurations on (or near) the Pareto frontier, optimally balancing model accuracy against deployment cost rather than optimizing solely for accuracy.
@@
Figure 14: End-to-end latency and accuracy trends under iteration budget constraints in ReAct. Markers indicate the points of maximum accuracy (red diamond) and peak cost-efficiency (blue diamond), as measured by accuracy-to-latency ratio.
@@
Figure 15: End-to-end latency and accuracy trends with varying numbers of few-shot examples in ReAct. Markers indicate the configuration with the highest accuracy (red diamond) and the peak cost-efficiency (blue diamond), based on normalized accuracy-to-latency ratio.
@@ V-B Test-Time Scaling of AI Agents
Figure 16: Accuracy-latency trade-offs with sequential, parallel scaling on HotpotQA. Legends denote the scaling level: maximum reflection steps in (a, b), and number of child nodes per expansion in (c).
@@
[…]
@@
Figure 17: Accuracy–cost trade-offs under test-time scaling across two model sizes (Llama-3.1-Instruct 8B and 70B) on HotpotQA. (a)–(c) compare Reflexion (top row) and LATS (bottom row) across latency, token usage, and energy consumption. While 70B achieves higher accuracy with fewer steps, the 8B model when paired with parallel scaling can approach 70B performance with lower energy cost. Each point denotes a different level of test-time reasoning.
@@
Figure 17(a) shows that both the 8B and 70B Llama-3.1-Instruct [llama_3.1_8b_instruct, llama_3.1_70b_instruct] models eventually reach saturation in accuracy, but they differ in how quickly they reach this point. The 70B model achieves high accuracy with relatively low latency, whereas the 8B model requires much longer inference times to reach similar performance. This trend is echoed in Figure 17(b), which plots total token usage. The 8B model consumes significantly more tokens at high-accuracy settings, indicating that it needs more reasoning iterations to match the 70B model’s performance. However, as shown in Figure 17(c), the 8B model is substantially more energy-efficient. While the 70B model relies on 8 A100 GPUs, the 8B model runs on just one, resulting in lower total energy consumption per request, even when requiring more reasoning steps to be involved.
Interestingly, the performance gap between models can be partially closed with effective scaling strategies. Reflexion (which uses sequential scaling) shows limited accuracy on the 8B model. But with LATS and parallel scaling, the 8B model achieves near-70B performance by exploring multiple paths and selecting the best one. This shows that a test-time strategy can play a compensatory role in low-resource settings.
VI AI Infrastructure Implications
In this section, we analyze the system-level impact of agentic test-time scaling by quantifying the GPU energy consumption and datacenter-wide power demands of AI agents relative to conventional single-turn LLM inference. Following the methodology in Section V-B, this section utilizes Reflexion and LATS as representative AI agents that employ sequential and parallel scaling, respectively. Reflexion and LATS design points were selected based on the highest-accuracy configurations in Figure 17. Llama-3.1-Instruct 8B and 70B models are used as backend LLMs and ShareGPT serves as the baseline for conventional single-turn inference.
@@ VI AI Infrastructure Implications
Based on recent estimates, ChatGPT serves roughly 500 million to 1.27 billion weekly active users (WAU) [chatgpt_wau_500M, similarweb_chatgpt_2025, demandsage_chatgpt_stats, forbes_chatgpt_1b_2025], which corresponds to approximately 71.4 million to 181.4 million daily active users (DAU). Assuming the conservative estimate of 71.4 million DAU and that each user submits just a “single” agentic query per day, Reflexion’s daily GPU energy consumption would be approximately 2.97 GWh for the 8B model and 24.89 GWh for the 70B model. Although our analysis does not account for LLM request batching [orca, vllm], which can amortize execution overheads, the estimate remains conservative for three reasons: (1) it represents a lower-bound based on the conservative DAU estimate of 71.4 million and assumes just one query per user, despite accelerating adoption and increasing user demand, (2) it includes only GPU energy, omitting overheads from CPU, memory, networking, storage, and cooling, and (3) even the larger 70B model considered in our study is orders of magnitude smaller than today’s large-scale LLMs, which now reach hundreds of billions to trillions of parameters [meta2025llama4, moonshot2025kimi, deepseekai2025deepseekr1, felloai2024gemini15].
@@
TABLE III: Accuracy, latency, and GPU energy consumption when servicing a single agent request on HotpotQA. Numbers in parentheses indicate the relative increase over ShareGPT (the conventional single-turn inference).
Accuracy (%) Latency (seconds) Energy (Wh/query)
8B ShareGPT 4.23 (1×) 0.32 (1×)
Reflexion 38 649.34 (153.7×) 41.53 (130.9×)
LATS 80 380.90 (90.1×) 22.76 (71.7×)
70B ShareGPT 6.40 (1×) 2.55 (1×)
Reflexion 67 720.00 (112.6×) 348.41 (136.5×)
LATS 82 305.67 (47.8×) 158.48 (62.1×)
TABLE IV: Datacenter-wide power demand under current and future traffic scenarios (71.4 Million and 13.7 Billion Queries/day), assuming the HotpotQA benchmark.
Power @ 71.4 Million Queries/day (Watts) Power @ 13.7 Billion Queries/day (Watts)
8B ShareGPT 1.0 M 182.7 M
Reflexion 123.6 M 23.7 G
LATS 67.7 M 13.0 G
70B ShareGPT 7.6 M 1.5 G
Reflexion 1.0 G 198.9 G
LATS 471.5 M 90.5 G
[…]
Sustainability challenges of agentic test-time scaling. Collectively, our findings show that AI agent performance does not scale proportionally with the associated compute, energy, and power costs. Once accuracy saturates, additional test-time scaling yields diminishing returns while imposing substantial system-level burdens. This cost inefficiency is not merely theoretical; it poses concrete constraints on real-world deployments. For instance, OpenAI’s Deep Research [deep_research], designed for complex multi-step reasoning, can take up to 30 minutes per request [deep_research_faq]. To keep infrastructure costs manageable, OpenAI limits usage to 25 runs every 30 days for ChatGPT Plus users [deep_research_faq]. These limits highlight the financial and computational challenges of sustaining AI systems that rely heavily on intensive test-time computation.
@@
VII Discussion
Future directions for sustainable AI agent serving. While the primary objective of our work is to raise awareness within the community about the broader system-level implications of deploying agents—particularly the immense infrastructural costs associated with them—we also highlight several promising directions that we believe will be critical for sustainably serving agentic systems. First, conventional model-level optimizations such as quantization [lin2024awq, xiao2023smoothquant, frantar2022gptq], distillation [hinton2015distilling, muralidharan2024compact, sreenivas2024llm, guo2025deepseek], sparse architectures [fedus2022switch, yuan2025native, liang2024mixture], and adaptive model routing [panda2025adaptive, ding2025best] will remain essential for reducing computational and memory demands. For example, constructing multi-agent systems that combine a heterogeneous mix of small and large language models (SLMs and LLMs), and dynamically selecting the appropriate model depending on the agent’s role and task significance (e.g., planning vs. acting), can substantially reduce both operational cost and latency—an approach also advocated by [belcak2025small]. Second, for AI agents that are not strictly latency-sensitive, carbon-aware computing [radovanovic2022carbon, li2023clover, gsteiger2024caribou] that migrates parts of the execution to compute instances incurring lower carbon intensity or electricity costs can provide both environmental and economic benefits. Finally, adaptive scaling strategies [snell2024scaling, qu2025optimizing, yang2025towards] that dynamically adjust compute resources based on task difficulty and importance enable agents to allocate GPU resources more efficiently, avoiding over-provisioning while maintaining quality of service. Together, these directions highlight exciting opportunities for improving the efficiency, sustainability, and scalability of serving AI agent workloads.
Agent serving under SLA constraints. The industry has only recently begun exploring agentic systems, so there are currently no well-established or widely accepted SLA standards. Consequently, we did not (and realistically could not) conduct our analysis with respect to a specific SLA target. The goal of this work is to characterize energy–efficiency trade-offs across diverse agent configurations, rather than to optimize for a fixed latency constraint. Considering the current direction of agentic system development, agents are generally allowed to spend additional computation time to achieve higher reasoning quality. Our analysis intentionally examines this behavior to highlight the inefficiency of unrestricted test-time scaling and to motivate more energy-aware SLA design in future agent deployments. A detailed exploration of efficient agent serving under SLA constraints is left as future work.
VIII Related Work
AI agent workflows. Recent advances in LLM-based AI agents have introduced diverse workflows that combine language-based reasoning with external tool use. Single-agent frameworks (e.g., ReAct [react], Reflexion [reflexion], LATS [lats], LLMCompiler [llmcompiler]) enhance decision-making through iterative reasoning, tool execution, and reflection. Multi-agent systems, such as CAMEL [camel] and AutoGen [autogen], further extend these capabilities by structuring task execution, communication, and coordinated behaviors among multiple agents. While these workflows substantially improve their capabilities and behavioral flexibility, their system-level implications remain underexplored. This work provides the first comprehensive analysis of representative AI agents, offering insights into the efficiency and scalability of agentic systems.
@@ VIII Related Work
AI agent interfaces for tool-augmented reasoning. In parallel with behavioral advancements in AI agents, recent efforts have focused on standardizing AI agent APIs and protocols to facilitate broader integration and deployment. OpenAI’s function-calling interface [openai_function_calling] defines a structured mechanism for API invocation, enabling agents to interact with tools in a verifiable and consistent manner. Anthropic’s Model-Context-Protocol (MCP) [mcp] further formalizes how agents manage context and interact with tools. Google’s Agent-to-Agent (A2A) protocol [agent2agent] complements these efforts by specifying a standard for multi-agent communication. Although these contributions primarily standardize the interfaces and protocols for agent interaction, our work takes an orthogonal system-level perspective, uncovering the AI infrastructural challenges posed by agentic workloads under test-time scaling.
@@
System-level optimization of AI agents. LLMCompiler [llmcompiler], Alto [alto], and Ayo [ayo] reduce inference latency by enabling pipelined and parallel execution across reasoning steps. Autellix [autellix] optimizes latency through queue-aware scheduling, while AI Metropolis [ai_metropolis] and Murakkab [murakkab] improve multi-agent coordination and resource isolation. While these works focus on optimizing specific components such as scheduling or execution flow, our study provides a broader characterization of infrastructural behaviors and efficiency trade-offs across diverse AI agents at scale.
LLM inference optimization techniques. The AI community has only recently begun exploring agents, making it both unclear and highly challenging to determine the most effective methodology for applying various LLM-focused optimization techniques to agentic systems. To maintain clarity and generality in our analysis, this paper focuses on fundamental and widely adopted LLM inference optimizations that are readily available in existing AI serving frameworks [vllm, zheng2024sglang]. A comprehensive exploration of all the latest LLM optimizations in the literature for agentic systems is beyond the scope of this paper, so we provide a summary of recent LLM inference optimizations and discuss their applicability to agents below.
In terms of KV cache management, hierarchical caching [cachedattention, infinigen, multiturn] and non-prefix KV cache reuse [cacheblend] approaches extend the naive prefix caching, enabling more efficient KV cache reuse. Token pruning [dejavu, keyformer, attentionsink] or KV cache compression [minicache, kvcachegen], or model architectural improvement like grouped-query attention [ainslie2023gqa] and multi-head attention [liu2024deepseek] reduces the memory footprint of the KV cache, which will be especially helpful for agent workloads with long contexts. Regarding decoding, Speculative decoding [Leviathan2023SpecDecode] predicts multiple candidate tokens and validates in parallel to reduce decoding latency. In agents, speculative decoding can potentially become effective as their reasoning often generates predictable schema patterns (e.g., JSON structures or function arguments), which will increase the acceptance rate of speculative branches and improve overall decoding throughput. Prefill-decode disaggregation [splitwise, distserve, dynamo] allows flexible and efficient resource allocation by decoupling the compute-intensive prefill phase from the memory-bound decode phase. For agents with long-context that incur substantial prefill computation load, disaggregation mitigates interference between prefill and decoding workloads, leading to more stable performance and improved overall efficiency.
IX Conclusion